A working student has to deliver a package to one of their clients.Before he can reach their client's office,he has to travel 3 km due east,then 3 km due south,and then 7 km due west.How far is their office from the clients office?
Answer by rapaljer(4671)

Traveling 3 km due east and then 7 km due west is a net of 4 km west, along with 3 km south.

4^2 + 3^2 = x^2
16+9=x^2
25=x^2
x=5 km total distance from their office.

Dr. Robert J. Rapalje